two cirlces with a scale factor of 3:2, the larger circle area is 81pi. How do I find out the smaller circle's area?
Answer by checkley77(12569)

3:2 or 81/x
3/2=81/x
3x=81*2
3x=162
x=162/3
x=54pi. answer for the area of the smaller circle.